OpenStudy (daryan):

Prove the cofunction identity using the addition and subtraction formulas. tan(pi/2 -u)= cot u

OpenStudy (xapproachesinfinity):

use the identities they asked you to use what's stopping you? tan(a-b)=sin(a-b)/cos(a-b) sin(a-b)=sinacosb-sinbcosa cos(a-b)=cosacosb+sinasinb go now and work on it

OpenStudy (daryan):

According to my teacher, I have to use the formula tana+tanb/1-tanatanb

OpenStudy (freckles):

\[\tan(a+b)=\frac{\tan(a)+\tan(b)}{1-\tan(a)\tan(b)}\] then replace a with pi/2 and replace b with -u

OpenStudy (daryan):

I did that

OpenStudy (freckles):

ok and what did you get?

OpenStudy (freckles):

the only problem is tan(pi/2) is not going to work out

OpenStudy (freckles):

so i don't know why the teacher suggested that

OpenStudy (freckles):

or said you had to use it

OpenStudy (daryan):

I can't seem to break it down

OpenStudy (freckles):

i'm saying you can't use that identity your teacher provided because it doesn't work

OpenStudy (freckles):

tan(pi/2) doesn't exist

OpenStudy (freckles):

what i would do is draw a pretty little triangle

OpenStudy (freckles):

and assign variables to each of the sides of that triangle (being a right triangle)

OpenStudy (freckles):

then I would bind tan(pi/2-u) and cot(u)

OpenStudy (freckles):

and see if they were the same ratio

OpenStudy (freckles):